The following is the cell phone usage for the month of November in the Adams family.

Father Mother Daughter Son
price per minute in dollars .40 1.30 1.10 .29
probability of that person using the phone .6 .3 .2 .2

This table represents an accurate probability distribution
Answer #1

Answer

No, this table does not represents an accurate probability distribution because we know that the maximum value of probability in a distribution is sum of all individual probability, which must be equal to 1.

In this case, the total probability is greater than 1

P(father)+P(mother)+P(daughter)+P(son) = 0.6+0.3+0.2+0.2 = 1.3

So, it is clear that 1.3 is greater than 1, which makes the inaccurate probability distribution for the given data in the table.

So, we can say that this table does not represents an accurate probability distribution.